OpenStudy (anonymous):

4x - y ( line ) 2y + x HELP PLEASE

OpenStudy (mrnood):

Please post the whole question - this is not complete . You need to post equations (with an = sign) to make a line.... 4x-y is not an equation....

OpenStudy (anonymous):

It's a dividing line.

OpenStudy (anonymous):

\[\frac{4x - y}{2y + x}\]

OpenStudy (anonymous):

Ya like that

OpenStudy (anonymous):

There is no way to simplify it more, it is in simplified form..

OpenStudy (anonymous):

The answers are -3 9 1 18

OpenStudy (anonymous):

Please post your complete question or else you take screenshot and upload it here.. :)

OpenStudy (mrnood):

there are 2 variables - you cannot have 1 number as an answer as ^^ water says - post the whole question please

OpenStudy (anonymous):

My bad x=3 an y=3

OpenStudy (anonymous):

Then just put the values in. Wherever you are seeing x, replace it with 3 and y with 3.. Like this: \[\frac{4(3) - (3)}{2(3) + 3} = ??\]

OpenStudy (anonymous):

So -3?

OpenStudy (anonymous):

How solve it properly and slowly..

OpenStudy (anonymous):

Solve numerator first; 4*3 = ??

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Good.. And 12 - 3 = ??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Okay Good.. SO, numerator = 9, keep it aside.. Now come to denominator: (2*3) + 3 = ??

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

So, finally you got: \[\frac{Numerator}{Denominator} = \frac{9}{9} = ??\]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

So, what option you are seeing??

OpenStudy (anonymous):

1 would be my answer

OpenStudy (anonymous):

Good, Well Done!!!

OpenStudy (anonymous):

Which option is correct, then??